### v3

# MedFuncta: A Unified Framework for Learning Efficient Medical Neural Fields

Abstract
Research in medical imaging primarily focuses on discrete data representations that poorly scale with grid resolution and fail to capture the often continuous nature of the underlying signal. Neural Fields (NFs) offer a powerful alternative by modeling data as continuous functions. While single-instance NFs have successfully been applied in medical contexts, extending them to large-scale medical datasets remains an open challenge. We therefore introduce MedFuncta, a unified framework for large-scale NF training on diverse medical signals. Building on Functa, our approach encodes data into a unified representation, namely a 1D latent vector, that modulates a shared, meta-learned NF, enabling generalization across a dataset. We revisit common design choices, introducing a non-constant frequency parameter $\omega$ in widely used SIREN activations, and establish a connection between this $\omega$-schedule and layer-wise learning rates, relating our findings to recent work in theoretical learning dynamics. We additionally introduce a scalable meta-learning strategy for shared network learning that employs sparse supervision during training, thereby reducing memory consumption and computational overhead while maintaining competitive performance. Finally, we evaluate MedFuncta across a diverse range of medical datasets and show how to solve relevant downstream tasks on our neural data representation. To promote further research in this direction, we release our code, model weights and the first large-scale dataset - MedNF - containing $500\text{\,}\mathrm{k}$\$ latent vectors for multi-instance medical NFs.
1 Introduction
It is a common choice to represent data on discretized grids, e.g., to represent an image as a grid of pixels. While this data representation is widely explored, it poorly scales with grid resolution and ignores the often continuous nature of the underlying signal [14]. Recent research demonstrated that NFs provide an interesting, continuous alternative to represent different kinds of data modalities like sound [52], images [53], shapes [38], videos [8], or 3D scenes [39], by treating data as neural functions that take spatial or temporal positions (e.g., pixel coordinates) as input and output the appropriate measurements (e.g., image intensity values).
Figure 1: Our proposed MedFuncta framework. A meta-learned shared network - with a novel $\omega$-schedule that acts as an implicit layer-wise learning rate scheduler - is conditioned by a signal-specific parameter vector to represent a medical signal. A signal can be reconstructed by evaluating the function $f_{\theta,\phi^{(i)}}$ over some spatial (or temporal) domain $\Omega$.
Single-instance NF training typically involves overfitting a neural network to a single signal [52, 53, 38, 8, 39]. While this training paradigm yields accurate representations for single instances, it is prohibitively expensive when scaled to large datasets. This scalability issue has gained particular importance as researchers increasingly explore using NFs as compressed dataset representations [15, 34, 47], where the neural network weights themselves are treated as a data modality. Naively training single-instance NFs additionally leads to highly unordered weight spaces across separately trained networks, which complicates downstream learning on the weights [49, 41]. Although specialized architectures designed to handle the permutation symmetries inherent to the multilayer perceptrons (MLPs) that typically comprise NFs exist [41, 49], alternative frameworks that avoid these issues in the first place can greatly simplify downstream learning. To overcome these challenges, this work introduces a framework, called MedFuncta, that generalizes medical NFs from isolated, single-instance models to dataset-level neural representations. The central idea, borrowed from Functa [14] and shown in Figure 1, is to meta-learn a shared neural representation across the dataset, in which each signal is represented by a unique, signal-specific parameter vector (also referred to as latent) that conditions a shared network. This structure enables the model to capture and reuse redundancies across different signals, drastically improving computational efficiency and scalability. Unlike prior methods that rely on patch-based representations [15, 5], our proposed framework represents each signal, from 1D time series to 3D volumetric data, with a single 1D latent vector. This abstraction enables consistent downstream processing across diverse data types, and is especially advantageous in medical applications, where the ability to unify multiple data modalities under a common representation is desirable [^1] Also see Appendix A., and where the inherent capability of NFs to handle irregularly sampled, heterogeneous data provides further benefits.
Our main contributions are threefold:
  1. 1.
    Optimization of Learning Dynamics at Scale: We revisit common design choices with respect to learning dynamics and generalization. In contrast to prior work that employs SIREN activations with a single frequency parameter $\omega$ across all layers, we demonstrate that introducing a non-constant, layer-dependent $\omega$-schedule significantly improves both convergence speed and reconstruction quality. We provide theoretical insights into the interplay between a layer’s $\omega$-parameter and its effective learning rate, connecting these results to recent research on theoretical learning dynamics. We further validate these findings through comprehensive ablation studies.
  2. 2.
    Efficient Meta-Learning with Context Reduction: To deal with the scalability issues of high-dimensional medical data, we propose an efficient meta-learning framework for shared network training. By utilizing sparse supervision, this context-reduced meta-learning framework allows us to significantly reduce memory consumption and computational overhead while maintaining competitive performance and speeding up the learning process.
  3. 3.
    Demonstration of the Utility of MedFuncta Across Datasets and Tasks: We extensively evaluate our framework across a diverse range of medical signals of various modalities and dimensions, and further demonstrate that we can solve relevant downstream tasks on our proposed representation. To accelerate community research, we open-source our implementation, trained network weights, and a comprehensive dataset - MedNF - containing $500\text{\,}\mathrm{k}$\$ latent vectors for multi-instance NFs in medical imaging and machine learning (similar to [48, 34], but for medical data).
2 Related Work
2.1 Neural Fields
Neural Fields [62] model signals as continuous neural functions that map from some domain to the corresponding signal value. Recent advances in NFs have mainly focused on alleviating the spectral bias [45] of commonly used MLPs. A variety of different strategies evolved, ranging from the application of Fourier Features in ReLU MLPs [57], over periodic sinusoidal activation functions like SIREN [52], to using Gabor wavelet-based nonlinearities like WIRE [46], or multi-resolution hash-grid encodings like in InstantNGP [40]. Research not only focused on network architectures and activation functions, but also explored training strategies like soft mining [27], or context-pruning [55] that both aim to identify informative samples, as well as network initialization strategies [66, 23]. In the medical field, single-instance NFs have become particularly relevant in settings where data is scarce, irregularly sampled [21], or should be free of cohort priors [35]. Applications include slice-to-volume and unsupervised dynamic MR reconstruction [63, 17], self-supervised motion correction [1], sparse-view CBCT reconstruction [67], and image registration [61, 50].
2.2 Generalizable Neural Fields
Figure 2: (Left) The proposed network with shared parameters $\theta$, that is conditioned by a single signal-specific parameter vector $\phi^{(i)}$. (Right) The proposed meta-learning strategy that, starting from a random initialization of $\theta$, learns shared network parameters $\theta^{*}$ in such a way that we can fit an unseen signal by only updating $\phi^{(i)}$ for very few steps.
Several approaches have been proposed to generalize NFs from single instances to entire datasets. Earlier methods like SIREN [52], GASP [13], COIN [12], or HyperDiffusion [16] rely on first fitting a set of single-instance NFs, and subsequently using the resulting weights for solving downstream tasks like generation or compression. This is not only expensive, but the unordered nature of the constructed weight spaces also limits the performance of these methods. To bypass first-stage, single-instance NF fitting, auto-decoder frameworks [43, 6, 54] jointly optimize a shared network and signal-specific latent vectors. Representative methods such as DeepSDF [43], CINA [10] and NISF [54] follow this strategy, but their naive joint optimization makes fitting new latents at test time slow and computationally expensive. To overcome this limitation, frameworks that learn-to-learn, i.e., train the shared network in a way that fitting a new latent at test time requires few steps only, arose. Methods like MetaSDF [51], COIN++ [15], Functa [14], SpatialFuncta [5] or LIFT [25] use different meta-learning strategies to achieve this goal. A parallel line of research aims to train hypernetworks that generate NF weights, without the need for first-stage single-instance NF fitting [11, 29, 69]. They do so by using the auto-decoder framework to learn hypernetworks that map from a low-dimensional, locally linear manifold to the desired NF weights [11], or by leveraging an explicit field parameterization and score field networks to learn distributions over neural fields in an end-to-end fashion [69].
While single-instance and generalizable NFs have been studied as a novel data modality in the deep learning community [14, 48, 34], their application to medical data, along with the unique challenges and opportunities, remains unexplored.
3 Method
In this work, we aim to find parameter-efficient NFs for $N$ signals $\big\{s_{1},...,s_{N}:\mathbb{R}^{C}\to\mathbb{R}^{D}\big\}$, e.g., a set of time series, or images, by learning a functional representation of the signal $s_{i}$ given some context set $\mathcal{C}^{(i)}:=\big\{(\mathbf{x}_{j},\mathbf{y}_{j})\big\}_{j=1}^{M}$ with $M$ coordinate-value pairs $(\mathbf{x}_{j},\mathbf{y}_{j})\in\mathbb{R}^{C}\times\mathbb{R}^{D}$. While parameterizing such a function as a neural network $f_{\theta}:\mathbb{R}^{C}\to\mathbb{R}^{D}$, with all parameters $\theta$ being optimized to fit a single signal is widely explored [39, 46, 52], this approach is prohibitively expensive when scaled to large datasets. As we aim to find an efficient framework, we want to reduce the number of parameters optimized per signal by reusing redundant information. Additionally, neural network weights exhibit permutation symmetries, i.e., reordering neurons within a network might change the weight matrix but not necessarily the represented function. This makes direct operations on the weights $\theta$ difficult. [41] As we want to use these NFs as a data representation that we can learn on, we require an alternative parameterization.
In the following section, we will therefore:
  • introduce a suitable architecture that effectively exploits redundancies in medical signals and introduce an $\omega$-schedule that improves the network’s learning dynamics (Section 3.1),
  • propose a scalable meta-learning approach for training this network, compensating for high-computational requirements by reducing the used context set, effectively training with sparse supervision (Section 3.2),
  • and demonstrate how to use our proposed approach at test time (Section 3.3).
3.1 Network Architecture
We argue that most sets of signals (datasets) contain large amounts of redundant information or structure that we can learn over the entire set. This is particularly true in medicine, where patients exhibit broadly similar yet slightly varying anatomies. We therefore define a neural network $f_{\theta,\phi^{(i)}}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{D}$ with shared network parameters $\theta$ that represent this redundant information and additional signal-specific parameters $\phi^{(i)}\in\mathbb{R}^{P}$ that condition the base network to represent a specific signal $s_{i}$. We apply a $K$-layer MLP architecture with a hidden dimension of $L$ and FiLM modulated SIREN activations [36, 52], where all layers $k\in\{2,...,K-1\}$ are defined as:
$x\mapsto\sin\Big(\omega_{k}\big(\underbrace{\mathbf{W}_{k}x+\mathbf{b}_{k}}_{\text{Linear}}+\underbrace{\mathbf{m}_{k}(\phi^{(i)})}_{\text{Modulation}}\big)\Big),$ (1)
with $\omega_{k}$ being the layer’s frequency parameter, $\mathbf{W}_{k}$ and $\mathbf{b}_{k}$ being the weights and biases of the $k$-th layer, and $\mathbf{m}_{k}(\cdot)$ being a linear layer that maps the signal-specific parameters $\phi^{(i)}$ to a shift-modulation vector that is added in the base network’s nonlinearity [44]. The first layer is a SIREN layer that projects the input coordinate to a higher-dimensional space. The last layer is a linear layer that performs a simple mapping to the desired output dimension. An overview of the proposed architecture is shown in Figure 2.
3.1.1 Network Initialization
A proper initialization of NFs has been shown to have a huge influence on convergence and overall performance of the applied networks [23, 66]. We therefore initialize the network’s weights and biases similar to Sitzmann et al. [52]:
$\mathbf{W}_{k},\mathbf{b}_{k}\sim\mathcal{U}\left(-\frac{\sqrt{6/n}}{\omega_{k}},\frac{\sqrt{6/n}}{\omega_{k}}\right),$ (2)
with $n$ being the layer’s input dimension. The first layer’s weights and biases are initialized as $\textbf{W}_{1},\textbf{b}_{1}\sim\mathcal{U}(-1/n,1/n)$}.
3.1.2 Introducing an $\boldsymbol{\omega}$-Schedule
While recent research treats $\omega$ as a single hyperparameter that remains constant over all network layers [52, 14], we identify this as a main restriction when being applied in a generalization setting. Following recent work demonstrating that shallow network layers learn the low-frequency content of a signal and deeper layers add more and more high-frequency information [7], we introduce an $\boldsymbol{\omega}$-schedule that linearly increases from $\omega_{1}$ to $\omega_{K}$. We find that this is equivalent to a layer-wise learning rate schedule, which positively influences the network’s learning dynamics.
3.1.3 Equivalence to Layer-wise lr-Schedule
To see this, we consider two layers with indices $m$ and $n$ and different $\omega$-values $\omega_{m}\neq\omega_{n}$. For the sake of simplicity, we treat all variables as scalars [^2] A comprehensive derivation of all equations and the extension to the general vectorial case can be found in Appendix B.:
$y$ $=\sin(\omega_{m}(\mathbf{W}_{m}x+\mathbf{b}_{m}))$ (3)
$y$ $=\sin(\omega_{n}(\mathbf{W}_{n}x+\mathbf{b}_{n})).$ (4)
We further consider the case where both layers are initialized with the same underlying random values. Specifically, let $\tilde{\mathbf{W}}^{(0)},\tilde{\mathbf{b}}^{(0)}\sim\mathcal{U}(-\sqrt{6/n},\sqrt{6/n})$ and define:
$\begin{split}\mathbf{W}_{m}^{(0)}&=\tfrac{1}{\omega_{m}}\tilde{\mathbf{W}}^{(0)},\quad\mathbf{b}_{m}^{(0)}=\tfrac{1}{\omega_{m}}\tilde{\mathbf{b}}^{(0)}\\ \mathbf{W}_{n}^{(0)}&=\tfrac{1}{\omega_{n}}\tilde{\mathbf{W}}^{(0)},\quad\mathbf{b}_{n}^{(0)}=\tfrac{1}{\omega_{n}}\tilde{\mathbf{b}}^{(0)}.\end{split}$ (5)
This ensures that both layers produce the same output upon initialization. We now examine the condition that needs to be fulfilled in order for the two layers to still produce the same output after a single update step, i.e., we want to find out when $\omega_{m}\mathbf{W}_{m}^{(1)}=\omega_{n}\mathbf{W}_{n}^{(1)}$ holds. The corresponding gradient descent updates for the layer weights are:
$\mathbf{W}_{m}^{(1)}$ $=\mathbf{W}_{m}^{(0)}-\tau_{m}\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial\mathbf{W}_{m}}$ (6)
$\mathbf{W}_{n}^{(1)}$ $=\mathbf{W}_{n}^{(0)}-\tau_{n}\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial\mathbf{W}_{n}},$ (7)
where $\tau_{m}$ and $\tau_{n}$ are the learning rates of the respective layers. By plugging in Equations (6) and (7), we get:
$\omega_{m}\underbrace{\left[\mathbf{W}_{m}^{(0)}-\tau_{m}\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial\mathbf{W}_{m}}\right]}_{\mathbf{W}_{m}^{(1)}}\overset{!}{=}\omega_{n}\underbrace{\left[\mathbf{W}_{n}^{(0)}-\tau_{n}\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial\mathbf{W}_{n}}\right]}_{\mathbf{W}_{n}^{(1)}}$.
Since we know that the initialization of both layers is equal, this equation reduces to:
$\omega_{m}\tau_{m}\frac{\partial y}{\partial\mathbf{W}_{m}}=\omega_{n}\tau_{n}\frac{\partial y}{\partial\mathbf{W}_{n}}.$ (9)
By simply plugging in the partial derivatives:
$\frac{\partial y}{\partial\mathbf{W}_{m}}$ $=\cos(\omega_{m}(\mathbf{W}_{m}x+\mathbf{b}_{m}))\omega_{m}x$ (10)
$\frac{\partial y}{\partial\mathbf{W}_{n}}$ $=\cos(\omega_{n}(\mathbf{W}_{n}x+\mathbf{b}_{n}))\omega_{n}x,$ (11)
and further simplifying the equation, we finally end up with the condition:
$\frac{\tau_{n}}{\tau_{m}}=\left(\frac{\omega_{m}}{\omega_{n}}\right)^{2}.$ (12)
This means that two layers with different $\omega$-values show the same behavior, if we rescale the learning rates according to the inverse quadratic relationship $\tau\propto\frac{1}{\omega^{2}}$. If we view this observation the other way around, we find that changing $\omega$ across different layers is similar to an implicit layer-wise learning rate schedule.
3.1.4 A Connection to the Network’s Learning Dynamics
This observation offers a so far overlooked perspective on the $\omega$-parameter in SIREN networks and establishes a connection to recent research on learning dynamics, providing a theoretical justification for introducing the proposed $\omega$-schedule. Chen et al. [7] showed that shallow MLP layers yield faster convergence due to more informative gradients and a smoother loss landscape. They formalize this as the layer convergence bias, arguing that training strategies that prioritize low-frequency representations in shallow layers, while deferring high-frequency details to deeper layers, achieve better performance. They further find that shallow layers tolerate higher learning rates, whereas deeper layers begin to effectively learn once the learning rate decays. Our perspective on the $\omega$-parameter in SIRENs naturally fits into this framework. By gradually increasing $\omega$ with depth, we effectively lower the learning rate of deeper layers. This enforces a staged optimization dynamic, where shallow layers first stabilize around smooth, low-frequency features, and deeper layers subsequently refine high-frequency details. We validate this theoretical insight through ablation studies (Section 4.6), demonstrating that networks incorporating our proposed $\omega$-schedule outperform current state-of-the-art networks with a constant $\omega$-parameter.
3.2 Efficient Meta-Learning with Context Reduction
Figure 3: (Left) The proposed approach for meta-learning the shared model parameters $\theta$. We perform $G$ inner-loop updates before performing a single meta-update. To compensate for the expensive second-order optimization, we reduce the context set $\mathcal{C}_{\mathtt{red}}$ in the inner-loop. An Algorithm describing the full meta-learning approach can be found in Appendix C. (Right) The proposed test time adaptation scheme. There is no need to update $\theta$ at test time, which allows us to make use of the full context set $\mathcal{C}$.
To efficiently create a set of NFs, we aim to meta-learn the shared parameters $\theta$ such that we can fit a signal $s_{i}$ by only optimizing $\phi^{(i)}$ for very few update steps [15, 14] (see Figure 2). We follow a CAVIA approach [70], shown in Figure 3, by defining an optimization process over the shared model parameters:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\mathtt{MSE}}(\phi_{G}^{(i)},\theta;\mathcal{C}^{(i)}),$ (13)
where in each meta/outer-loop update step (i.e., the optimization of $\theta$), the inner-loop optimizes $\phi^{(i)}$ from scratch ($\phi_{0}^{(i)}:=\mathbf{0}$), performing $G$ update steps:
$\phi_{g+1}^{(i)}:=\phi_{g}^{(i)}-\alpha\nabla_{\phi}\mathcal{L}_{\mathtt{MSE}}(\phi_{g}^{(i)},\theta,\mathcal{C}^{(i)})$ (14)
using stochastic gradient descent (SGD) with a fixed learning rate $\alpha$. The meta-update is performed using AdamW [33] with a learning rate $\beta$ that follows a cosine annealing learning rate schedule [32]. All optimization steps aim to minimize the reconstruction error when evaluating the learned function $f_{\theta,\phi^{(i)}_{g}}$ on a given context set $\mathcal{C}^{(i)}$, by minimizing the mean squared-error (MSE) loss:
$\mathcal{L}_{\mathtt{MSE}}(\phi^{(i)}_{g},\theta;\mathcal{C}^{(i)}):=\frac{1}{|\mathcal{C}^{(i)}|}\sum_{j\in\mathcal{C}^{(i)}}\|f_{\theta,\phi^{(i)}_{g}}(\mathbf{x}_{j})-\mathbf{y}_{j}\|_{2}^{2}$.
Performing a single meta-update step involves backpropagating through the entire inner-loop optimization, which requires retaining the computational graph in GPU memory to compute second-order gradients [18]. [^3] Updating $\theta$ necessitates backpropagation through all inner-loop parameters $\phi_{1:G}$, each of which is itself a function of $\theta$. Consequently, computing the update for $\theta$ involves Hessian-vector products, which in turn demand storing the complete inner-loop computational graph. More information can be found in Appendix C. This resource-intensive task does not scale well to high-dimensional signals. While first-order approximations [18, 42] or auto-decoder training approaches that do not rely on second-order optimization exist [43], recent research has shown that this results in severe performance drops or unstable training [14, 15]. To overcome this limitation, we propose to make use of a reduced context set $\mathcal{C}_{\mathtt{red}}^{(i)}$ during the inner-loop optimization [55]. This reduced context set contains a subset of the full context set $\mathcal{C}_{\mathtt{red}}^{(i)}\leq\mathcal{C}^{(i)}$, thus saving GPU memory that is required for second-order optimization. We obtain the reduced context set by randomly sampling $\gamma|\mathcal{C}^{(i)}|$ coordinate-value pairs from $\mathcal{C}^{(i)}$. We empirically find that reducing the selection ratio $\gamma$ results in marginal performance drops, while significantly reducing the required GPU memory and speeding up the training (see Table 5).
3.3 Fitting Neural Fields at Test Time
Given the meta-learned model parameters $\theta^{*}$, we fit a NF to each signal $s_{1},...,s_{N}$, by optimizing the signal-specific parameter vectors $\phi^{(1)},...,\phi^{(N)}$. We start with initializing a signal-specific parameter vector $\phi^{(i)}:=\mathbf{0}$ and optimize $\phi^{(i)}$ for $H$ steps by minimizing $\mathcal{L}_{\mathtt{MSE}}(\phi^{(i)},\theta^{*};\mathcal{C}^{(i)})$. We do this for all $N$ signals. As no second-order optimization is required at test time (see Figure 3), we can make use of the full context set $\mathcal{C}^{(i)}$, i.e., we use all the available information at test time. A set of NFs representing the signals $s_{1},...,s_{N}$ is therefore defined by the network architecture, the shared model parameters $\theta^{*}$, and the signal-specific parameters $\phi^{(1)},...,\phi^{(N)}$. While meta-learning $\theta^{*}$ requires solving a complex optimization problem, fitting a NF at test time (i.e., optimizing $\phi^{(i)}$) simply requires $H$ SGD updates, which results in fast and low-resource inference ($0.5\text{\,}\mathrm{s}$\$ and $1\text{\,}\mathrm{GB}$\$ GPU memory for a $64\times 64$ image), a desirable property in medical applications.
4 Experiments
4.1 Datasets
We conduct experiments on a diverse set of publicly available datasets, spanning medical signals of different modalities, to demonstrate the flexibility of our proposed method: a single-lead ECG dataset [22], two different chest X-ray datasets [26, 60], a retinal OCT dataset [26], a fundus image dataset [30], a dermatoscopy image dataset [9, 58], a colon histopathology dataset [24], a cell microscopy dataset [31], a brain MRI dataset [3, 4, 37], and a lung CT dataset [2]. No preprocessing was needed for the ECG dataset. For all 2D datasets, we use preprocessed versions from MedMNIST [64, 65]. The brain MRI and lung CT datasets were preprocessed as described in [19].
4.2 Implementation Details
All networks were trained with $G=10$ inner-loop and $H=20$ test time adaptation steps. The inner-loop learning rate was set to $\alpha=10^{-2}$ and the outer-loop learning rate to $\beta=3\times 10^{-6}$. Network configurations and further training details are reported in Table 1. All experiments were carried out on a single NVIDIA A100 ($40\text{\,}\mathrm{GB}$) GPU. Our code is publicly available at https://github.com/pfriedri/medfuncta.
Table 1: The number of layers $K$, the hidden dimension $L$, the batch size $B$, the context selection ratio $\gamma$, $\omega_{1}$ and $\omega_{K}$, as well as the representation size $P$.
Signal Dim. $K$ $L$ $B$ $\gamma$ $\omega_{1}$ $\omega_{K}$ $P$
1D 88 $64$ $64$ $1.00$ 2020 $200$ $64$
2D $15$ $256$ $24$ $0.25$ 2020 $400$ $2048$
3D $15$ $256$ 44 $0.25$ 2020 $300$ $8192$
4.3 Reconstruction Quality
We first validate that our proposed approach can fit a wide range of medical signals by performing reconstruction experiments. We meta-learn the shared network parameters $\theta$ on a training set and evaluate the reconstruction quality on a hold-out test set. All models are trained for a fixed number of $250\text{\,}\mathrm{k}$ iterations, and testing is performed using the weights that achieve the best validation scores.
Table 2: Mean reconstruction quality of our proposed method, evaluated on a hold-out test set after meta-learning for $250\text{\,}\mathrm{k}$ iterations. MSE scores are multiplied by $10^{3}$. The spatial dimensions are 1D: $187$, 2D: $64\times 64$, 3D: $32\times 32\times 32$.
Dim. Signal Type MSE $(\downarrow)$ PSNR $(\uparrow)$ SSIM $(\uparrow)$ LPIPS $(\downarrow)$
1D ECG [22] $0.086$ $43.301$ $0.964$
2D Chest X-ray [60] $0.097$ $40.719$ $0.985$ $0.013$
Chest X-ray [26] $0.146$ $39.301$ $0.977$ $0.014$
Retinal OCT [26] $0.203$ $37.321$ $0.934$ $0.071$
Fundus Camera [30] $0.054$ $43.151$ $0.978$ $0.006$
Dermatoscope [9, 58] $0.133$ $40.273$ $0.962$ $0.023$
Colon Histopathology [24] $0.943$ $31.886$ $0.925$ $0.021$
Cell Microscopy [31] $0.013$ $49.944$ $0.994$ $0.008$
3D Brain MRI [3, 4, 37] $0.130$ $39.191$ $0.993$
Lung CT [2] $1.561$ $28.325$ $0.913$
We measure mean squared error (MSE), peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS) [68] and report the results in Table 2. Qualitative examples of the performed reconstruction experiments are shown in Figure 4.
InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRecon.Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 4: Input and reconstruction examples from the hold-out test set for (from left to right) chest X-ray [60], chest X-ray [26], retinal OCT [26], fundus camera [30], dermatoscope [9, 58], colon histopathology [24], and cell microscopy [31] images.
While performance is generally better on homogeneous datasets, where redundancies can more effectively be exploited, the proposed method also learns to represent complex inhomogeneous datasets, such as the colon histopathology dataset. [^4] Additional qualitative results can be found in Appendix G.
4.4 Scaling MedFuncta to High-Resolution Signals
To highlight our proposed approach’s computational efficiency and scalability, we additionally evaluate its performance on higher-resolution signals. We, therefore, perform additional reconstruction experiments over multiple datasets, using images with a resolution of $128\times 128$ and $224\times 224$ as supervision signals. Reconstruction scores after $250\text{\,}\mathrm{k}$ meta-learning steps are reported in Table 3.
Table 3: Mean reconstruction quality of MedFuncta on higher resolutions. We use the setup from Section 4.2, only changing batch size $B$, representation size $P$, selection ratio $\gamma$, $\omega_{1}=30$, and $\omega_{K}=300$ . We also report the required training GPU memory in GB. MSE scores are multiplied by $10^{3}$.
Dataset Resolution $B$ $P$ $\gamma$ MSE $(\downarrow)$ PSNR $(\uparrow)$ SSIM $(\uparrow)$ Mem. $(\downarrow)$
Chest X-ray [60] $128\times 128$ 88 $8192$ $0.25$ $0.216$ $37.174$ $0.952$ $28.68$
$224\times 224$ 44 $16384$ $0.10$ $0.401$ $34.510$ $0.909$ $25.39$
Dermatoscope [9, 58] $128\times 128$ 88 $8192$ $0.25$ $0.277$ $37.072$ $0.906$ $28.68$
$224\times 224$ 44 $16384$ $0.10$ $0.472$ $34.752$ $0.920$ $25.39$
Qualitative results are shown in Figure 5. The results demonstrate that our proposed method can reconstruct high-resolution signals, even when being trained on a single $40\text{\,}\mathrm{GB}$ GPU only. We believe that larger networks and longer, distributed training would further improve performance, especially on the $224\times 224$ data, which runs under substantially different conditions due to hardware constraints.4
[…]
Figure 5: Input and reconstruction examples from the hold-out test set for chest X-ray [60], and dermatoscope [9, 58] images.
4.5 Classification Experiments
To assess whether the learned representation captures relevant information about the underlying signal, we perform classification experiments on the signal-specific parameters $\phi$ [14, 41], using a $k$-Nearest-Neighbor ($k$-NN) classifier, or a 3-layer MLP with ReLU activations and dropout. We compare these simple classifiers on our MedFuncta representation to ResNet50 [20] and EfficientNet-B0 [56] on the original data, and report the number of network parameters, training time, accuracy, and F1 scores. All models were trained for 50 epochs using AdamW with a learning rate of $10^{-3}$. The scores in Table 4 show the classification performance on a hold-out test set based on the model parameters yielding the highest validation accuracy.
Table 4: Classification Performance. We report the number of network parameters, the training time in seconds, accuracy, as well as F1 scores.
Dataset (Classes) Classifier Param. Time $(\downarrow)$ Acc. $(\uparrow)$ F1 $(\uparrow)$
Chest X-ray [26] (2) $k$-NN $(k=1)$ on $\phi$ 0 0 $81.57$ $0.87$
$k$-NN $(k=3)$ on $\phi$ 0 0 $80.93$ $0.87$
MLP on $\phi$ $1.2\times 10^{6}$ $45$ $\mathbf{89.10}$ $\mathbf{0.88}$
ResNet50 $23.5\times 10^{6}$ $450$ $83.49$ $0.80$
EfficientNet-B0 $4.0\times 10^{6}$ $270$ $84.46$ $0.82$
Dermatoscope [9, 58] (7) $k$-NN $(k=1)$ on $\phi$ 0 0 $68.98$ $0.38$
$k$-NN $(k=3)$ on $\phi$ 0 0 $69.28$ $0.32$
MLP on $\phi$ $1.2\times 10^{6}$ $65$ $\mathbf{74.96}$ $0.48$
ResNet50 $23.5\times 10^{6}$ $700$ $74.36$ $\mathbf{0.49}$
EfficientNet-B0 $4.0\times 10^{6}$ $410$ $70.42$ $0.44$
We find that solving the two classification tasks (binary and multi-class) on our proposed representation $\phi$ generally works well. We outperform both ResNet50 and EfficientNet-B0, applied to the original images, in terms of accuracy and can demonstrate competitive F1 scores, while requiring less training time and model parameters. These results indicate that our proposed representation actually captures informative features of the underlying signals. The observed performance improvement may be attributable to removing redundant signal components, which are represented by $\theta$ and not $\phi$.
4.6 Ablation Studies
[…]
Table 5: The effect of the context selection ratio $\gamma$ on the reconstruction quality and GPU memory required for training. Measured on chest X-ray dataset [60] $(64\times 64)$ after $100\text{\,}\mathrm{k}$ iterations and a batch size of 12, using the baseline configuration. We also report the training time for 100 iterations, averaged over $50\text{\,}\mathrm{k}$ iterations following $2\text{\,}\mathrm{k}$ warm-up iterations.
Selection Ratio $(\gamma)$ $0.1$ $0.25$ $0.5$ $0.75$ $1.0$
PSNR (dB) $34.33$ $35.62$ $36.22$ $36.53$ $36.60$
SSIM $0.941$ $0.955$ $0.958$ $0.959$ $0.959$
Memory [GB] $6.77$ $11.43$ $19.32$ $28.66$ $34.34$
Time [s] / 100 iterations $37.38$ $42.04$ $68.75$ $83.18$ $91.61$
To assess the effectiveness of our proposed $\boldsymbol{\omega}$-schedule, we sweep over a combination of different configurations with $w_{1}=\{10,20,30,40,50\}$ and $\omega_{K}:=\delta\omega_{1}$, with $\delta=\{1,2,5,10,20\}$. The results, shown in Figure 6, demonstrate that applying our proposed schedule consistently improves the performance over setups with a single $\omega$-parameter across all layers. Among the tested configurations, $\omega_{1}=20$ and $\omega_{K}=400$ demonstrated the most favorable results. We further observe a performance drop for very large $\omega_{K}$-values, which can be attributed to training collapse. Reducing $\omega_{K}$ has proven effective in mitigating such instability, especially for high-resolution or high-dimensional signals.
(a) PSNR
(b) SSIM
Figure 6: Grid search over different $\omega_{1}$ and $\omega_{K}:=\delta\omega_{1}$ parameters. We report (a) PSNR, and (b) SSIM values after 25k iterations. Outliers with red borders were excluded from color scaling to ensure enough contrast in the relevant regions. All experiments were conducted on the chest X-ray dataset [60] $(64\times 64)$.
Table 6: The effect of introducing a global learning rate schedule, our proposed $\omega$-schedule, or a reduced context set $\mathcal{C}_{\mathtt{red}}$ with $\gamma=0.25$, as well as a combination of all of them in comparison to Functa [14]. Measured on chest X-ray dataset [60] $(64\times 64)$ after $250\text{\,}\mathrm{k}$ training iteratons. MSE scores are multiplied by $10^{3}$.
Method BS Param. Memory (GB) MSE $(\downarrow)$ PSNR $(\uparrow)$ SSIM $(\uparrow)$ LPIPS $(\downarrow)$
Functa [14] $12$ $17.1\times 10^{6}$ $25.13$ $0.403$ $34.304$ $0.940$ $0.059$
Ours (Baseline) $12$ $7.7\times 10^{6}$ $34.34$ $0.179$ $37.836$ $0.970$ $0.016$
Ours  + Global lr-sched $12$ $7.7\times 10^{6}$ $34.34$ $0.168$ $38.282$ $0.973$ $0.017$
Ours  + $\omega$-Schedule $12$ $7.7\times 10^{6}$ $34.34$ $0.119$ $39.684$ $0.979$ $0.013$
Ours  + $\mathcal{C}_{\mathtt{red}}$$(\gamma=0.25)$ $24$ $7.7\times 10^{6}$ $21.51$ $0.205$ $37.338$ $0.968$ $0.019$
Ours  + All $24$ $7.7\times 10^{6}$ $21.51$ $\mathbf{0.097}$ $\mathbf{40.719}$ $\mathbf{0.985}$ $\mathbf{0.013}$
As a last experiment, we compare our baseline approach [^5] Configuration of our baseline: $K=15$, $L=256$, $\gamma=1.0$, $\omega=30$, $G=10$, $H=20$, no global learning rate schedule. to Functa [^6] Configuration of Funtca[14]: $K=15$, $L=512$, $\gamma=1.0$, $\omega=30$, $G=3$, $H=3$, no global learning rate schedule.[14] and study the effect of introducing a global learning rate schedule, our proposed $\omega$-schedule, or the proposed context reduction scheme $\mathcal{C}_{\mathtt{red}}$ on the reconstruction performance. The results in Table 6 show that our approach outperforms Functa by a PSNR of $\sim$6.4\text{\,}\mathrm{dB}$\$ and that each of the proposed components consistently improves upon the baseline. While the context reduction scheme does not enhance reconstruction performance, the substantial GPU memory savings outweigh the minor decrease in reconstruction quality, which we consider a practical improvement. Furthermore, we observe that training with reduced context and an increased batch size leads to more stable training. Contrary to Functa [14], we find that applying a smaller network, but allocating more GPU resources to a longer inner-loop optimization can be beneficial. Figure 7 shows a qualitative comparison of our proposed approach MedFuncta (Ours + All) and Functa. [^7] Additional comparisons can be found in Appendix F.
InputReconstructionDifferenceReconstructionDifferenceMedFuncta (Ours)Functa [14]Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Qualitative comparison between our proposed approach MedFuncta and Functa [14], with absolute difference maps.
5 Dataset: MedNF
While neural fields, or neural network weights in general, emerged as a novel data modality in the deep learning community, large-scale datasets like Implicit-Zoo [34] or Model-Zoo [48] are scarce and not available for the medical domain. To promote research on weight space learning, finding well-performing architectures that operate on our representation, or solving further downstream tasks, we release a large-scale dataset, called MedNF [^8] A detailed description of the datasets can be found in Appendix D., that contains seven sub-datasets - ChestNF, PathNF, DermaNF, OctNF, PneumoniaNF, RetinaNF, TissueNF - with a total of more than $500\text{\,}\mathrm{k}$ NFs. Each of these NFs consists of a network shared across a sub-dataset and a signal-specific parameter vector $\phi$. Depending on the dataset, we provide labels for binary-/multi-class classification or ordinary regression problems. We release this dataset at https://doi.org/10.5281/zenodo.14898708.
6 Limitations & Future Work
While our proposed approach demonstrates promising results, we identify several limitations and directions for future research. First, the expressiveness of current MLP-based NFs remains a key bottleneck, as even relatively large models struggle to capture high-dimensional, complex signals. Developing more expressive architectures is therefore a promising research direction, particularly in the largely overlooked context of generalization. Second, the conditioning mechanism used in this paper incurs a substantial parameter overhead. Future work could explore parameter-sharing strategies or alternative conditioning schemes like signal-specific low-rank adaptations of weights and biases. Third, we believe that our novel view on SIRENs $\omega$-parameter and its relation to the network’s learning dynamics should be further explored. While substantial progress has been made on NF architectures, activations, and initialization, the role of NF learning dynamics is an interesting yet underexplored direction. Finally, numerous downstream tasks like inpainting, registration, or generation are worth being exploring within our framework. We release the MedNF dataset to support research in this direction, aiming to facilitate the development of efficient, high-performing architectures on our proposed representation and encourage the community to investigate downstream applications.
7 Discussion
Motivated by recent progress in learning generalizable NFs, we present MedFuncta, a unified framework for efficiently learning large-scale neural representations of medical signals. We not only introduce a novel $\omega$-schedule for commonly used SIRENs, effectively enforcing a staged optimization process through layer-wise learning rates, we also address the computational overhead of previous methods by presenting an efficient meta-learning framework that utilizes training with sparse supervision. We validate our findings through extensive experiments and ablations across a wide range of diverse medical signals and demonstrate that we can solve relevant downstream tasks on our proposed neural representation.
Acknowledgments
This work was financially supported by the Werner Siemens Foundation through the MIRACLE II project. JM is supported by Bavarian State Ministry for Science and Art (Collaborative Bilateral Research Program Bavaria – Québec: AI in medicine, grant F.4-V0134.K5.1/86/34).
Appendix A Why We Don’t Rely on Patch-Based Representations
Although there is a substantial body of work on generalizable patch-based neural representations, such as COIN++ [15], SpatialFuncta [5], and LIFT [25], the focus of this work is different. We aim to learn a shared representation across signals of varying modalities and dimensions, where each signal is encoded as a single 1D latent vector. While we recognize that the inductive bias of patch-based representations can be advantageous, and while we believe that our findings could also be utilized within these frameworks, such an exploration lies beyond the scope of this work. We argue that a unified representation across diverse signal types is particularly valuable in medicine, where integrating data from multiple modalities remains an active yet unsolved research challenge.
Appendix B Relation Between $\boldsymbol{\omega}$-Schedule and Layer-Wise Learning-Rates
In this section, we provide a comprehensive mathematical derivation of Section 3.1.3 and extend it to the general vectorial case.
B.1 Problem Formulation
Consider two SIREN layers with different frequency parameters $\omega_{m}\neq\omega_{n}$. The layers are defined :
$\mathbf{y}_{m}$ $=\sin(\omega_{m}(\mathbf{W}_{m}\mathbf{x}+\mathbf{b}_{m}))$ (16)
$\mathbf{y}_{n}$ $=\sin(\omega_{n}(\mathbf{W}_{n}\mathbf{x}+\mathbf{b}_{n})),$ (17)
where $\mathbf{W}_{m},\mathbf{W}_{n}\in\mathbb{R}^{p\times d}$ are weight matrices, $\mathbf{b}_{m},\mathbf{b}_{n}\in\mathbb{R}^{p}$ are bias vectors, $\mathbf{x}\in\mathbb{R}^{d}$ is an input vector, and $\mathbf{y}_{m},\mathbf{y}_{n}\in\mathbb{R}^{p}$ are output vectors. Our goal is to determine the relationship between the learning rates $\tau_{m}$ and $\tau_{n}$ that ensures both layers maintain equivalent outputs throughout training, despite having different frequency parameters.
B.2 Initialization Analysis
To understand how different $\omega$-values affect a layer’s learning rate, we consider the case where both layers are initialized with the same underlying random values but different scaling factors (as described in Section 3.1.1). Specifically, let $\tilde{\mathbf{W}}^{(0)},\tilde{\mathbf{b}}^{(0)}\sim\mathcal{U}(-\sqrt{6/n},\sqrt{6/n})$ and define:
$\mathbf{W}_{m}^{(0)}$ $=\frac{1}{\omega_{m}}\tilde{\mathbf{W}}^{(0)},\quad\mathbf{b}_{m}^{(0)}=\frac{1}{\omega_{m}}\tilde{\mathbf{b}}^{(0)}$ (18)
$\mathbf{W}_{n}^{(0)}$ $=\frac{1}{\omega_{n}}\tilde{\mathbf{W}}^{(0)},\quad\mathbf{b}_{n}^{(0)}=\frac{1}{\omega_{n}}\tilde{\mathbf{b}}^{(0)}.$ (19)
This ensures that:
$\omega_{m}\mathbf{W}_{m}^{(0)}$ $=\omega_{n}\mathbf{W}_{n}^{(0)}=\tilde{\mathbf{W}}^{(0)}$ (20)
$\omega_{m}\mathbf{b}_{m}^{(0)}$ $=\omega_{n}\mathbf{b}_{n}^{(0)}=\tilde{\mathbf{b}}^{(0)},$ (21)
which means that both layers produce the same output upon initialization:
$\sin\big(\omega_{m}(\mathbf{W}_{m}^{(0)}\mathbf{x}+\mathbf{b}_{m}^{(0)})\big)=\sin\big(\omega_{n}(\mathbf{W}_{n}^{(0)}\mathbf{x}+\mathbf{b}_{n}^{(0)})\big).$
B.3 Identifying the Learning Rate Relation
We now examine the condition that needs to be fulfilled for both layers to maintain equal behavior after gradient descent updates. We therefore require:
$\omega_{m}\mathbf{W}_{m}^{(1)}\overset{!}{=}\omega_{n}\mathbf{W}_{n}^{(1)}.$ (23)
The gradient descent update steps are defined as:
$\mathbf{W}_{m}^{(1)}$ $=\mathbf{W}_{m}^{(0)}-\tau_{m}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{m}}$ (24)
$\mathbf{W}_{n}^{(1)}$ $=\mathbf{W}_{n}^{(0)}-\tau_{n}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{n}},$ (25)
where $\tau_{m}$ and $\tau_{n}$ are the learning rates for layers $m$ and $n$, respectively. Substituting the update equations (24) and (25) into condition (23), we get:
$\omega_{m}\left[\mathbf{W}_{m}^{(0)}-\tau_{m}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{m}}\right]=\omega_{n}\left[\mathbf{W}_{n}^{(0)}-\tau_{n}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{n}}\right].$
Using the initialization condition (20), this simplifies to:
$\omega_{m}\tau_{m}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{m}}=\omega_{n}\tau_{n}\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{n}}.$ (27)
For a given loss function $\mathcal{L}$, the gradients with respect to the weight matrices are defined as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{m}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{m}}\frac{\partial\mathbf{y}_{m}}{\partial\mathbf{W}_{m}}$ (28)
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{n}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{n}}\frac{\partial\mathbf{y}_{n}}{\partial\mathbf{W}_{n}},$ (29)
with:
$\frac{\partial\mathbf{y}}{\partial\mathbf{W}_{m}}$ $=\omega_{m}\mathbf{x}^{\top}\otimes\mathrm{diag}\Big(\cos\big(\omega_{m}(\mathbf{W}_{m}\mathbf{x}+\mathbf{b}_{m})\big)\Big)$ (30)
$\frac{\partial\mathbf{y}}{\partial\mathbf{W}_{n}}$ $=\omega_{n}\mathbf{x}^{\top}\otimes\mathrm{diag}\Big(\cos\big(\omega_{n}(\mathbf{W}_{n}\mathbf{x}+\mathbf{b}_{n})\big)\Big),$ (31)
where $\otimes$ is the Kronecker product and $\mathrm{diag}(\cdot)$ is the diagonal matrix with the entries of its argument on the diagonal. Substituting the gradient expressions from (30) and (31), Equation (27) becomes:
$\begin{aligned} &\omega_{m}\tau_{m}\Bigg(\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{m}}\,\omega_{m}\mathbf{x}^{\top}\otimes\mathrm{diag}\Big(\cos\big(\omega_{m}(\mathbf{W}_{m}\mathbf{x}+\mathbf{b}_{m})\big)\Big)\Bigg)\\ &\quad=\omega_{n}\tau_{n}\Bigg(\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{n}}\,\omega_{n}\mathbf{x}^{\top}\otimes\mathrm{diag}\Big(\cos\big(\omega_{n}(\mathbf{W}_{n}\mathbf{x}+\mathbf{b}_{n})\big)\Big)\Bigg)\end{aligned}$
Due to the initialization conditions (20) and (21), the arguments of the cosine functions are the same:
$\omega_{m}(\mathbf{W}_{m}^{(0)}\mathbf{x}+\mathbf{b}_{m}^{(0)})=\omega_{n}(\mathbf{W}_{n}^{(0)}\mathbf{x}+\mathbf{b}_{n}^{(0)}).$ (33)
The same holds true for the loss gradients with respect to the model outputs - considering the similar output from Equation (22):
$\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{m}}=\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{n}}$ (34)
Equation (32) therefore reduces to:
$\omega_{m}^{2}\tau_{m}=\omega_{n}^{2}\tau_{n},$ (35)
which can be reformulated as:
$\frac{\tau_{n}}{\tau_{m}}=\left(\frac{\omega_{m}}{\omega_{n}}\right)^{2}$ (36)
Appendix C Additional Details on the Meta-Learning Approach
This section provides further insights into our meta-learning framework, presented in Section 3.2 and summarized in Algorithm 1. We start with the inner-loop optimization process in which the signal-specific parameter vectors are initialized as a zero-vector:
$\phi^{(i)}_{0}:=\mathbf{0},$ (37)
and updated for g=0,,G1g=0,...,G-1 update steps using stochastic gradient descent (SGD):
$\phi^{(i)}_{g+1}=\phi^{(i)}_{g}-\alpha\nabla_{\phi}\mathcal{L}_{\mathtt{MSE}}\big(\phi^{(i)}_{g},\theta;\mathcal{C}^{(i)}\big).$ (38)
After performing $G$ inner-loop update steps, the meta-objective, used for updating the shared network parameters $\theta$, is defined as:
$\mathcal{L}_{\text{meta}}=\frac{1}{B}\sum_{i=1}^{B}\mathcal{L}_{\mathtt{MSE}}\big(\phi^{(i)}_{G},\theta;\mathcal{C}^{(i)}\big),$ (39)
for a batch with $B$ signals, where $\phi^{(i)}_{G}$ denotes the adapted signal-specific parameters, which are themselves functions of $\theta$. The gradient of the meta-objective with respect to $\theta$ is therefore defined as:
[…]
Computing the meta-gradient requires differentiating through the inner-loop optimization process. Each signal-specific parameter vector $\phi^{(i)}_{G}$ is the result of $G$ gradient descent steps that depend on the current shared parameters $\theta$. Therefore, when taking the derivative of the meta-objective with respect to $\theta$, we must account for both:
  • Direct effect: the explicit influence of $\theta$ on the loss $\mathcal{L}_{\mathtt{MSE}}$ given the adapted parameters $\phi^{(i)}_{G}$.
  • Indirect effect: the influence of $\theta$ on the loss through its effect on the inner-loop parameters $\phi^{(i)}_{G}$.
While taking the direct effect into account is straightforward, the indirect effect is captured by the term:
$\left(\frac{\partial\phi^{(i)}_{G}}{\partial\theta}\right)^{\top}\nabla_{\phi}\mathcal{L}_{\mathtt{MSE}}(\phi^{(i)}_{G},\theta;\mathcal{C}^{(i)}),$ (41)
which involves second-order derivatives of the inner-loop loss. To compute this term, we differentiate through the inner-loop recursion, which yields the following recursive formula for the Jacobian of the adapted parameters with respect to $\theta$:
$\frac{\partial\phi^{(i)}_{g+1}}{\partial\theta}=\frac{\partial\phi^{(i)}_{g}}{\partial\theta}-\alpha\left(\frac{\partial^{2}\mathcal{L}_{\mathtt{MSE}}(\phi^{(i)}_{g},\theta;\mathcal{C}^{(i)})}{\partial\theta\,\partial\phi}+\\ \frac{\partial^{2}\mathcal{L}_{\mathtt{MSE}}(\phi^{(i)}_{g},\theta;\mathcal{C}^{(i)})}{\partial\phi^{2}}\frac{\partial\phi^{(i)}_{g}}{\partial\theta}\right).$
This recursion explicitly shows how the inner-loop updates propagate the influence of $\theta$ to the adapted parameters, and why second-order terms (Hessian-vector products) are required to compute the full meta-gradient. After computing the required gradient, we take a single meta-update of the shared parameters $\theta$ using AdamW with learning rate $\beta$:
$\theta\;\leftarrow\;\texttt{AdamW}\big(\theta,\nabla_{\theta}\mathcal{L}_{\text{meta}},\beta\big).$ (43)
Algorithm 1 Meta-Learning NFs with Context Reduction During Inner-Loop Optimization
1:Dataset $\mathcal{D}=\{s_{1},s_{2},...,s_{N}\}$, inner-loop steps $G$, inner-loop learning rate $\alpha$, meta learning rate $\beta$, selection ratio $\gamma$, batch size $B$
2:while not converged do
3:  $\{s_{i}\}_{i=1}^{B}\sim\mathcal{D}$ $\triangleright$ Sample batch of $B$ signals
4:  $\mathcal{L}_{\text{meta}}\leftarrow 0$ $\triangleright$ Reset meta-loss
5:  for each signal $s_{i}$ in batch do
6:   $\phi_{0}^{(i)}=\mathbf{0}$ $\triangleright$ Initialize $\phi^{(i)}$
7:   for $g=0$ to G1G-1 do $\triangleright$ Inner-loop optim.
8:     $\mathcal{C}_{\mathtt{red}}^{(i)}\sim\mathtt{Sample}\big(\mathcal{C}^{(i)},\,\gamma\,|\mathcal{C}^{(i)}|\big)$
9:     $\phi_{g+1}^{(i)}\leftarrow\phi_{g}^{(i)}-\alpha\nabla_{\phi}\mathcal{L}_{\mathtt{MSE}}(\phi_{g}^{(i)},\theta;\mathcal{C}_{\mathtt{red}}^{(i)})$
10:   end for
11:   $\mathcal{L}_{\text{meta}}\leftarrow\mathcal{L}_{\text{meta}}+\mathcal{L}_{\mathtt{MSE}}(\phi_{G}^{(i)},\theta;\mathcal{C}^{(i)})$
12:  end for
13:  $\mathcal{L}_{\text{meta}}\leftarrow\frac{1}{B}\mathcal{L}_{\text{meta}}$
14:  $\theta\leftarrow\texttt{AdamW}\big(\theta,\nabla_{\theta}\mathcal{L}_{\text{meta}},\beta\big)$ $\triangleright$ Meta-update
15:  Update $\beta$ according to learning rate schedule
16:end while
17:return $\theta$
Appendix D Details on MedNF
Table 7: A list of all released sub-datasets with their original data source, their modality, the task we provide labels for, the number of samples, as well as the training/validation/test split. All datasets were adapted from MedMNIST [64, 65].
Dataset Orig. Source Modality Task Samples Train / Val / Test License
ChestNF [60] Chest X-Ray Multi-Label (14) Binary-Class (2) Classification 112 120112\,120 78 46878\,468 / 11 21911\,219 / 22 43322\,433 CC BY 4.0
PathNF [24] Colon Pathology Multi-Class (9) Classification 107 180107\,180 89 99689\,996 / 10 00510\,005 / $7180$ CC BY 4.0
DermaNF [9, 58] Dermatoscope Multi-Class (7) Classification 10 01510\,015 $7007$ / $1003$ / $2005$ CC BY-NC 4.0
OctNF [26] Retinal OCT Multi-Class (4) Classification 109 309109\,309 97 47797\,477 / 10 83210\,832 / $1000$ CC BY 4.0
PneumoniaNF [26] Chest X-Ray Binary-Class (2) Classification $5856$ $4708$ / $524$ / $624$ CC BY 4.0
RetinaNF [30] Fundus Camera Ordinal Regression (5) $1600$ $1080$ / $120$ / $400$ CC BY 4.0
TissueNF [31] Kidney Cortex Microscope Multi-Class (8) Classification 236 386236\,386 165 466165\,466 / 23 64023\,640 / 47 28047\,280 CC BY 4.0
In this section, we will provide further information on our MedNF datasets. All datasets, listed in Table 7, are adapted from MedMNIST [64, 65]. We meta-learned every model on the respective training set for $250\text{\,}\mathrm{k}$ iterations, using the setup described in Section 4.2. All datasets were derived from images with a resolution of $64\times 64$. We also aim to release models trained on higher and mixed resolutions in the future.
ChestNF is build upon the NIH-ChestXray14 [60] dataset and contains 112 120112\,120 frontal-view chest X-Ray images. It also provides binary disease-labels for the following diseases: (0) atelectasis, (1) cardiomegaly, (2) effusion, (3) infiltration, (4) mass, (5) nodule, (6) pneumonia, (7) pneumothorax, (8) consolidation, (9) edema, (10) emphysema, (11) fibrosis, (12) pleural, (13) hernia.
PathNF is build upon the NCT-CRC-HE-100K [24] dataset and contains 107 180107\,180 non-overlapping image patches of hematoxylin and eosin-stained colorectal cancer histology slides. It also provides a class label for one of nine tissues: (0) adipose, (1) background, (2) debris, (3) lymphocytes, (4) mucus, (5) smooth muscle, (6) normal colon mucosa, (7) cancer-associated stroma, (8) colorectal adenocarcinoma epithelium.
DermaNF is build upon the HAM10000 [9, 58] dataset and contains 10 01510\,015 dermatoscopic images of common pigmented skin lesions. It also provides a class label for one of the seven following diseases: (0) actinic keratoses and intraepithelial carcinoma, (1) basal cell carcinoma, (2) benign keratosis-like lesions, (3) dermatofibroma, (4) melanoma, (5) melanocytic nevi, (6) vascular lesions.
OctNF is build upon a dataset from [26] and contains 109 309109\,309 optical coherence tomography (OCT) images for retinal diseases. It also provides a class label for one of the four following diseases: (0) choroidal neovascularization, (1) diabetic macular edema, (2) drusen, (3) normal.
PneumoniaNF is build upon a dataset from [26] and contains $5856$ pediatric chest X-Ray scans. It contains the following binary label: (0) normal, (1) pneumonia.
RetinaNF is build upon the DeepDRiD [30] dataset and contains $1600$ retina fundus images. It provides labels for a 5-level grading of diabetic retinopathy severity.
TissueNF is build upon the BBBC051 [31] dataset from the Broad Bioimaging Benchmark Collection and contains 236 386236\,386 human kidney cortex cell images. It provides labels for one of the eight following cell types: (0) collecting duct, connecting tubule, (1) distal convoluted tubule, (2) glomerular endothelial cells, (3) interstitial endothelial cells, (4) leukocytes, (5) podocytes, (6) proximal tubule segments, (7) thick ascending limb.
Appendix E Negative Results
While the following explorations did not lead to promising results, we include them for completeness and transparency. We believe that reporting negative results helps to clarify the scope of our contributions, reduce redundant future efforts, and provide insight into design choices that, while theoretically appealing, did not prove effective in practice.
As our method constructs a representation space, i.e., the space in which the signal-specific parameter vectors $\phi$ reside, we explored incorporating a supervised contrastive loss [28] into our meta-learning objective. While this led to modest improvements in a simple binary classification task, we observed no improvement in more challenging multi-class classification settings. We assume that this is due to the lack of a learned encoder, usually available in representation learning settings. Consequently, we did not pursue this approach further.
We also experimented with quantized representations, where each entry is drawn from a learned codebook. Inspired by ideas from vector-quantized autoencoders [59], we hypothesized that this could improve performance by simplifying the decoding task for the shared network. However, we observed no such benefits and therefore retained continuous representations.
Based on the observation that shallow layers primarily capture relatively simple, low-frequency features while deeper layers encode more complex, high-frequency details, we hypothesized that a pyramid-like network structure might improve performance [7]. In this design, shallow layers would contain fewer neurons, reflecting the lower complexity of low-frequency features, while deeper layers would be allocated more neurons to better model the harder high-frequency details. However, this approach did not yield substantial improvements.
Lastly, instead of randomly sampling a subset of pixels in our context reduction approach, we experimented with gradient-based context pruning, as proposed in [55]. However, this method did not outperform random subsampling and, in some cases, led to unstable and collapsing training. We therefore decided to continue using random subsampling.
Appendix F Additional Comparisons to Functa
In this section, we present additional comparisons to Functa [14]. In Figure 8, we plot the evolution of PSNR, MSE, SSIM, and LPIPS on the validation set over the course of training. Our method not only achieves superior scores in significantly less time, it also exhibits more stable training dynamics. We also provide additional qualitative comparisons to Functa, along with absolute difference maps, in Figure 9. As in Figure 7, all models were trained following the protocol described in Section 4.6. Results are shown for our best-performing setup (”Ours + AllAll”). The presented training curves and images are from the same runs.
(a) PSNR
(b) MSE
(c) SSIM
(d) LPIPS
Figure 8: Development of PSNR, MSE, SSIM, and LPIPS on the validation set throughout a training run ($250\text{\,}\mathrm{k}$ iterations).
InputReconstructionDifferenceReconstructionDifferenceMedFuncta (Ours)Functa [14]Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Qualitative comparison between our proposed approach MedFuncta and Functa [14], with absolute difference maps.
Appendix G Additional Qualitative Results
In this section, we provide further qualitative results for the experiments in Section 4.3 and Section 4.4. The results are shown in Figure 10 - Figure 20.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Additional qualitative results on chest X-Ray [60] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 11: Additional qualitative results on chest X-Ray [60] images $(128\times 128)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 12: Additional qualitative results on chest X-Ray [60] images $(224\times 224)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 13: Additional qualitative results on chest X-Ray [26] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 14: Additional qualitative results on retinal OCT [26] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 15: Additional qualitative results on fundus camera [30] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 16: Additional qualitative results on dermatoscope [9, 58] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 17: Additional qualitative results on dermatoscope [9, 58] images $(128\times 128)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 18: Additional qualitative results on dermatoscope [9, 58] images $(224\times 224)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 19: Additional qualitative results on histopathology [24] images $(64\times 64)$.
Recon.InputRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 20: Additional qualitative results on cell microscopy [31] images $(64\times 64)$.